OpenStudy (tester97):

There are 15 students in a social studies class. Two students will be selected to present their term projects today. In how many different orders can two students be selected?

OpenStudy (valpey):

15 could be picked first times the 14 who could be picked to go second.

OpenStudy (anonymous):

15c2

OpenStudy (anonymous):

take example of 3 students and 2 should be picked, combinations are (1,2),(1,3),(2,3) which is 3c2 same goes here

OpenStudy (tester97):

im still not understanding it........

OpenStudy (anonymous):

Order seems to matter yet repetition of the same student does not so we can say it is a permutation with repetition and the formula for that is\[n^r\] where n is the number of students and r is the number of students we choose

OpenStudy (tester97):

so it would be? \[15^{2}\]

OpenStudy (anonymous):

Let students= 1,2,3,....15 combinations are (1,2),(1,3),(1,4)....(1,15) =========14 sets (2,3),(2,4).............(2,15)==========13 sets (3,4),(3,5)...........(3,15)===========12 sets “ “ .............................. ..............................(14,15)=======================1 set total= 14+13+12+11+10+9+8+7+6+5+4+3+2+1=105

OpenStudy (valpey):

Right. It doesn't seem to matter the order so divide my answer by two.

OpenStudy (luigi0210):

Use the formula: \(\Large C=\frac{n!}{r!(n-r)!}\)

OpenStudy (anonymous):

If order does not matter then we have a combination of things where repetition and order do not matter and the formula for that is\[\frac{ (n+r-1)! }{ r!(n-1)! }\]

OpenStudy (anonymous):

N is the number of students and r is the number of students we choose

OpenStudy (kropot72):

There are 15 choices for the first student, and each of these 15 can be combined with 14 choices for the second student. This gives 15 * 14 = 210 ways of choosing the 2 students where the order of selection matters.

OpenStudy (tester97):

Thanks guys i think i got it now ;)

OpenStudy (valpey):

15 students can be picked to go first times 14 who could be picked to go second which is 15*14 permutations. But since order doesn't matter (we say that John goes first and Luigi goes second is the same combination as Luigi goes first and John second then we divide the permutations (15*14) by 2 to get 15*7.

OpenStudy (valpey):

But the last sentence does seem to suggest order matters

OpenStudy (kropot72):

Yes, effectively the question says order of selection is important.

OpenStudy (tester97):

wait now im confused :\ is it 105 different ways? or is it 210?

OpenStudy (kropot72):

Permutations are needed, so 210 ways is correct.

OpenStudy (tester97):

ok thanks

OpenStudy (kropot72):

You;re welcome :)